The population of Boomtown is currently 12000 and expected to grow by 3.5% over the next year. What will its population be by then?
The population of Dullsville, on the other hand, is currently 10000 and expected to decrease by 5% over the next year. What will its population be by then?

Answer by ikleyn(52776)

.
(a) The population of Boomtown is currently 12000 and expected to grow by 3.5%
over the next year. What will its population be by then?

(b) The population of Dullsville, on the other hand, is currently 10000 and expected
to decrease by 5% over the next year. What will its population be by then?
~~~~~~~~~~~~~~~~~~

(a)  12000*(1+0.035) = 12000*1.035 = 12420.      ANSWER


(b)  10000*(1-0.05)  = 10000*0.95  =  9500.      ANSWER

Solved, with explanations.